OpenStudy (anonymous):

If a = 1+2+3+4+ ... +n, and b = 1+2^3+3^3+...+n^3 Find n, if gcf(a,b)+lcm(a,b)=4422

ganeshie8 (ganeshie8):

use sum of first n numbers, sum of cubes formulas

OpenStudy (anonymous):

b has the curious property of being a squared. a is equal to \[(n^2+n)/2\]because there are n values averaging (n+1)/2, and n times that gives the equation above. b should therefore equal \[(n^4+2n^3+n^2)/4\]We can prove this by showing the relation for 1 and for n+1: it is clearly true for one, and for (n+1) we have: \[b=\sum_{1}^{n}n^3= ((n+1)^4+2(n+1)^3+(n+1)^2)/4=\]\[(n^4+4n^3+6n^2+4n+1+2(n^3+3n^2+3n+1)+n^2+2n+1)/4=\]\[(n^4+2n^3+n^2)/4+(4n^3+12n^2+12n+4)/4\] Notice that our original term (n^3, expressed according to our first formula) remains the same while on the right we have \[n^3+3n^2+3n+1=(n+1)^3\]Therefore b=a^2 Q.E.D. From that you have a very simple polynomial (think about what happens with the gcf and lcd of a and a^2) to solve.

OpenStudy (anonymous):

Ok, now i know that b=a^2 and a = (n^2+n)/2 So that means that gcf(a,b) = a and lcm(a,b)=a^2 then a+a^2 = 4422 a^2+a-4422=0 (a-66)(a+67) so a = 66, because it's possitive and (n^2+n)/2=66 n^2+n-132 = 0 (n-11)(n+12)=0 Therefore n = 11 !!! Thanks.

OpenStudy (anonymous):

Nice!

OpenStudy (anonymous):

Oh, you should find the value of b as well, just for fun. I'd do it as a sum of cubes, but if you want to be boring you can take 66^2.

OpenStudy (anonymous):

It's fun to watch b equaling a^2 all of the way up.

OpenStudy (anonymous):

I better memorize that formulas , because the demonstration may be fun, but seems laborious.

OpenStudy (anonymous):

There is also something about Pascal's triangle and sums of powers of n...